# “vcd2df” - Leveraging Data Science Insights for Hardware Security Research thanks: This work was supported by NSF Award 171858

### v1

@@ V Related Work
Early computational methods to assess the correctness and security of hardware designs often directly examined trace data for known patterns, such as one-hot encoding [18, 14]. Other researchers applied data mining [6, 19, 22] or temporal logic [9, 7, 8, 12] techniques. Specific the to security context, researchers have manually identified properties [20, 4, 5] and used software tools incorporating clustering and classification for RISC [34] and CISC [13] designs. Recent work has discovered subsets of hyperproperties [25, 11], which are defined as relationships over multiple traces. In this work, we use Spark to evaluate multiple traces of a single design (PicoRV32) to discover grammatically similar hyperproperties to these works.
@@ VI Conclusion
We have proposed and demonstrated “vcd2df”, a proof-of-concept package for studying hardware designs with existing data science techniques. We have distributed our package through the relevant package manages for Python and the R Project, and have shown some simple examples of how easily insights may be learned from data frames containing trace data.
@@ Acknowledgements
We thank the NSF for funding. We thank the CRAN network for their helpful comments while developing “vcd2df” for the R Project. We thank Andres Meza for the AKER testbench. We thank Intel Corporation, the Kastner Research Group, and the HWSec@UNC research group for helpful discussions. We thank our undergraduate researchers, Kendall Leonard and Hannah Pahama, for suggesting library support for interfacing with VCD files during their research experience in the summer of 2024.

### v3

@@ V Related Work
Python VCD Interfaces
There are existing interfaces from Python to VCD files, namely PyVCD [^22] https://pypi.org/project/vcdvcd/ and vcdvcd [^23] https://pypi.org/project/pyvcd/. Versus our tool, PyVCD only writes VCD files, and vcdvcd only supports ASCII formatted output, which scales poorly to moderately sized designs. We elected to convert directly to binary encoding for scalability of write and query operations.
@@
Early computational methods to assess the correctness and security of hardware designs often directly examined trace data for known patterns, such as one-hot encoding [13, 14]. Other researchers applied data mining [15, 16, 17] or temporal logic [18, 19, 20, 21] techniques. Specific the to security context, researchers have manually identified properties [22, 23, 24] and used software tools incorporating clustering and classification for RISC [25] and CISC [26] designs. Recent work has discovered subsets of hyperproperties [27, 10], which are defined as relationships over multiple traces. We use Spark to evaluate multiple traces of a single design (PicoRV32) to discover grammatically similar hyperproperties to these works.
@@ VI Conclusion
We have proposed and demonstrated “vcd2df”, a proof-of-concept package for studying hardware designs with existing data science techniques. We have distributed our package through the relevant package managers for Python and the R Project, and have shown simple examples of insights from data frames containing trace data. Research is ongoing to study transient execution CPU vulnerabilities, and initial results suggest improved expressibility versus the state-of-the-art specification miners.
@@ Acknowledgements
We thank the NSF for funding. We thank the CRAN network for help developing “vcd2df” for the R Project. We thank Andres Meza for the AKER testbench. We thank Intel Corporation, the Kastner Research Group, and the HWSec@UNC research group for helpful discussions. We thank Kendall Leonard and Hannah Pahama for suggesting library support for a VCD dataframe in their summer REU of 2024.